OpenStudy (anonymous):

write the product as a sum 7cos6xcos7x

OpenStudy (raden):

first, 7cos6xcos7x = 7/2 * (2cos6xcos7x) to simplify that in bracket, use the formula : 2cosAcosB = cos(A+B) + cos(A-B)

OpenStudy (anonymous):

how do you know what a and b equal?\

OpenStudy (raden):

let A=6x and B=7x

OpenStudy (anonymous):

so i got cos(6x+7x)+cos(6x-7x)

OpenStudy (anonymous):

Can you use the the Cosine Sum/Difference identity? \[\cos(a + b) = \cos a \cos b - \sin a \sin b\]I'll give you a hint. First multiply both sides by 7. Then if we let a = 6x and b = 7x, how could you re-arrange the identity to solve just for 7cos(6x)cos(7x)? @NorthGeorgiaBoy17

OpenStudy (raden):

dont forget about 7/2 infront ...

OpenStudy (anonymous):

I'm using a different identity.

OpenStudy (raden):

us should use 2 formulas to get a new formula :) cos(a+b)=cosacosb−sinasinb cos(a-b)=cosacosb+sinasinb --------------------------- + cos(a+b) + cos(a-b) = 2cosacosb @genius12

OpenStudy (anonymous):

hehehe =)

OpenStudy (raden):

:)

OpenStudy (raden):

btw, i have to go now....

OpenStudy (anonymous):

byebye.

OpenStudy (anonymous):

okay so i plugged 6x for A and 7x for B. Do i just simplify it out or do i leave it??

OpenStudy (raden):

bye ^ (infinity) :)

OpenStudy (anonymous):

lol =)

OpenStudy (anonymous):

@NorthGeorgiaBoy17 Like @RadEn mentioned, factor out 7/2. Then change 2cos(7x)cos(6x) and write it as a sum using the identity: 2cosAcosB = cos(A + B) + cos(A - B) --> Here our A = 7x and b = 6x. Just plug in and solve and then multiply whatever you get by 7/2 which we factored out earlier.

OpenStudy (anonymous):

im sorry but how can you factor out a 7/2?